OpenStudy (anonymous):

Karen wrote the following steps to simplify the expression: 4 – (2x + 3x – 1). Step Mathematical Statement Justification 0 4 – (2x + 3x – 1) Given 1 4 – 2x – 3x + 1 Distributive Property 2 4 + 1 – 2x – 3x Associative Property of Addition 3 (4 + 1) + (–2x – 3x) Associative Property of Addition 4 5 – 5x Combine Like Terms Which step of justification is incorrect, and what should the justification for that step be to simplify the expression?

OpenStudy (anonymous):

The answers: Step 4; Distributive Property Step 2; Commutative Property of Addition Step 3; Associative Property of Multiplication Step 1; Multiplication Property of Equality

OpenStudy (anonymous):

sorta same concept

OpenStudy (anonymous):

I think its the last choice, am I wrong?

OpenStudy (anonymous):

what do you think it is? remember what i said before... the key is knowing what all these properties are...

OpenStudy (anonymous):

what is the difference between the Associative Property of Addition and the Commutative Property of Addition?

OpenStudy (anonymous):

associative is grouping and commutative is order.... for example: associative property: a + (b + c) = (a + b) + c commutative property: a + b = b + a

OpenStudy (anonymous):

So wouldn't the answer be the second choice then?

OpenStudy (anonymous):

Because step 2 is in the wrong format and the right format would be Commutative Property of Addition since it is in order.

OpenStudy (anonymous):

hmm.. that's what i was thinking the answer is (the second choice)... but i'm having second thoughts....

OpenStudy (anonymous):

Why the second step is wrong in the way it is put.

OpenStudy (anonymous):

since everything else is in order, then yeah, i'd say the second choice is your answer...

OpenStudy (anonymous):

Well, thank you so much for your time and knowledge, you learn something new everyday day. Have a great day!

OpenStudy (anonymous):

yeah.. it's order.... second choice is your answer...

OpenStudy (anonymous):

Now onto learning parallel line proofs lol so much to do tonight...

OpenStudy (anonymous):

thank you for your great questions.. you have been a great student... and also made me rethink about those properties....:)

OpenStudy (anonymous):

The way my class teaches the math is hard for me to understand, you might be seeing some more of my posts if i don't understand them.

OpenStudy (anonymous):

no problem... (maybe..) :)

OpenStudy (anonymous):

If I do have to do more posts, I hope you're still online :D